OpenStudy (anonymous):

Which of the following represents the vertical asymptotes of the function f(x)= x/ x^2 - 2x - 8 ?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

A. x = -8 and x =1 B. x= -1 and x = 8 C. x = -4 and x = 2 D. x = -2 and x = 4 Those are my choices.

OpenStudy (michele_laino):

vertical asymptotes are positioned at values of \(x\), such that the denominator is equal to zero, so, we have to solve this quadratic equation: \(\huge x^2-2x-8=0\)

OpenStudy (anonymous):

How do I do that?

OpenStudy (michele_laino):

do you know how to solve a quadratic equation?

OpenStudy (michele_laino):

in that case, we can solve such equation, in another way, namely, you have to search, among your options, two numbers, such that their product is equal to \(-8\), and their sum is equal to \(2\)

OpenStudy (anonymous):

Like out of my choices?

OpenStudy (michele_laino):

for example, first option gives, -8, and 1, now: sum = -8+1= -7 which is different from 2, so option A, is a wrong option

OpenStudy (anonymous):

I say C.

OpenStudy (anonymous):

D*

OpenStudy (michele_laino):

I think C is wrong since we have -4, and 2, and: sum = -4+2=-2 which is different from 2

OpenStudy (michele_laino):

option D is the correct option! :)